# AirIO: Learning Inertial Odometry with Enhanced IMU Feature Observability air-io.github.io

### v1

@@ IV-C System Performance
For the seen sequences, as shown in I, RoNIN and TLIO fail to capture the aggressive drone maneuvers as they are designed for pedestrian navigation with clear motion patterns. AirIO significantly outperforms these baselines, improving by 56.2% in ATE and 62.8% in RTE over TLIO. Furthermore, without integrating external thrust information like IMO, AirIO exceeds IMO by 43.9% in ATE and 61.4% in RTE relying solely on IMU measurement.
Surprisingly, we found that AirIO demonstrates remarkable generalizability to the unseen sequences. For example, as shown in 8, while other methods experience serious drift, AirIO maintains precise estimating. This performance gap arises because these methods tend to overfit training data, excelling within the training distribution but failing to generalize to UNSEEN data outside this distribution. Compared to IMO, AirIO shows a significant improvement of 86.6% in ATE and 84.7% in RTE.
@@
EuRoC Dataset In Table. II, both RoNIN and TLIO show suboptimal performance compared to AirIO. After integrating EKF, AirIO further improves the accuracy, showing an average additional 17.4% improvement in ATE. Specifically, it achieves an improvement of 52.9% and 54.4% in ATE relative to RoNIN and TLIO, respectively. This success is largely attributed to effective uncertainty modeling. AirIO fully leverages the uncertainties learned from both the IMU kinematic model and the motion network, ensuring optimal data fusion and improved performance.
@@ V Ablation Study
To evaluate the impact of different feature representations, we conduct ablation studies on the real-world dataset EuRoC and the simulation dataset Pegasus. We conduct experiments with the following input representation variants:
@@
    Body: The network processes body-frame IMU measurements defined in Equation (1), as input signal.
@@
    Global: The network processes global-frame IMU as input. The global-frame IMU is transformed from raw IMU data, as defined in Equation (2).
@@
    Body + Attitude: Building upon i. Body, the network encodes drone orientation as auxiliary inputs.
@@
    Global + Attitude: Building upon ii. Global, the network encodes drone orientation as auxiliary inputs.
@@
    Body – $\mathbf{R}^{T}_{i}\mathbf{g}^{\mathcal{G}}$: Building upon i. Body, we remove the rotation-coupled gravity term $\mathbf{R}^{T}_{i}\mathbf{g}^{\mathcal{G}}$ from Equation (1).
@@
    Global – $\mathbf{g}^{\mathcal{G}}$: Building upon ii. Global, we remove the gravity term $\mathbf{g}^{\mathcal{G}}$ from Equation (2).
To ensure a fair comparison, i. Body, iii. Body + Attitude, and v. Body – $\mathbf{R}^{T}_{i}\mathbf{g}^{\mathcal{G}}$ leverage ground-truth orientation to transform estimated velocity into global frame.
The results of the ablation study are presented in Table. III. You may find more results in Appendix. A-E. From these results, we address the following questions:
@@ V-A How Effective is the Body-frame Representation?
Compared to ii. Global, i. Body demonstrates significantly higher precision on the Pegasus. This single step - switching from global frame to body frame - results in an average ATE improvement of 73.0%. i. Body on the real-world EuRoC dataset exhibits a similar trend, achieving a 53.1% improvement, as illustrated in Fig. 10.
We further evaluate ii. Global and i. Body using the Accuracy Area Under the Curve (AUC) metric, which quantifies the performance across various error thresholds. Specifically, we calculate the relative position error over a 5-second interval. As shown in 9, on both datasets, i. Body achieves higher accuracy across all error thresholds and reaches near-perfect accuracy much more quickly.
These results highlight how the body frame effectively captures implicit attitude information in IMU data. Simply encoding the body frame IMU features helps the network learn the drone’s dynamics. Moreover, we find that combining the body-frame representation with the attitude information iii. Body + Attitude is better than i. Body.
Figure 10: Performance of MH_04_difficult in EuRoC dataset across different representations. ii. Global (Left) is suboptimal. Simply switching IMU data to i. Body (Center) improves significantly. After encoding drone rotation iii. Body + Attitude (Right), it has further improvement.
V-B How Effective is the Attitude Encoding?
Encoding the drone’s attitude information enhances the network’s ability to capture the dynamics of drones. As illustrated in Fig. 10, on the EuRoC dataset, iii. Body + Attitude achieves a further 30.3% improvement than the i. Body, outperforming ii. Global by 67.3% in ATE.
Additionally, we evaluate attitude encoding on global-frame representation iv. Global + Attitude. This also yields an improvement of approximately 29.6% over ii. Global, validating the efficiency of encoding attitude information.
V-C Shall We Keep Gravity in the Feature Representation?
Many learning-based IO methods remove gravity from the IMU encoding [28]. However, as discussed in Section. III-A, gravitational acceleration is a critical component for body-frame representation as it explicitly embeds drone rotation through the term $\mathbf{R}^{T}_{i}\mathbf{g}^{\mathcal{G}}$, as shown in Equation 1.
After removing this rotation-coupled gravity term, i. Body will discard essential attitude information. Table. III demonstrates that v. Body –$\mathbf{R}^{T}_{i}\mathbf{g}^{\mathcal{G}}$ causes significant performance degradation: ATE increases by 90.0% on EuRoC and 98.7% on Pegasus dataset.
@@ V-C Shall We Keep Gravity in the Feature Representation?
Seq. Global (2) $-\mathbf{g}^{\mathcal{G}}$ Global Body (1) $-\mathbf{R}_{i}^{T}\mathbf{g}^{\mathcal{G}}$ Global $+$Attitude Body Body $+$Attitude
EuRoC
MH02 18.048 17.892 16.136 9.522 4.537 2.531
@@
Pegasus
TEST_1 22.115 9.689 9.534 6.151 5.167 3.418
TEST_2 15.745 15.185 14.729 5.373 4.862 5.719
TEST_3 19.467 26.960 3.580 14.545 3.982 1.786
Avg. 19.109 17.278 9.281 8.690 4.670 3.641
@@ VI Conclusion & Discussion
In this work, we address the challenges identified in the discussion of IMO[6] without extra information like thrust signal. We demonstrate that simply preserving body-frame IMU data, inclusive of gravitational acceleration, significantly improves IO performance on UAVs. Building upon this analysis, we propose the AirIO system, which further enhances accuracy by explicitly encoding attitude information and integrating it with an EKF. Notably, compared to IMO, our model generalizes effectively to trajectories not present in the training datasets and operates without relying on additional inputs such as thrust. We believe that the proposed work has broader implications for reliable state estimation in agile drone flight.
@@ Appendix A Appendix
A-A BlackBird Dataset
@@
A-A1 Seen and Unseen sequences separation
@@
A-A2 Training and Testing Sequences Separation
@@
TABLE IV: Separation of trajectory sequences into SEEN and UNSEEN categories, and their respective allocations to training, validation, and testing sets.
@@
A-B Qualitative Evaluation for Blackbird dataset
We present more details on the evaluation of the Blackbird dataset. As shown in 11, we showcase seven additional trajectories that further highlight our method’s performance. Our method achieves superior performance in the SEEN sequences. For more than half of these sequences, it outperforms existing methods that rely on additional information in addition to IMU measurements, while our method uses only IMU data. When evaluating UNSEEN sequences, our model outperforms all existing methods on all sequences, demonstrating its remarkable adaptability.
(a) SEEN Sequences. From top to bottom: Clover, Egg, Winter and Star sequence. Our method demonstrates robust performance without requiring any additional sensor information.
(b) UNSEEN Sequences. From top to bottom: bentDice, Oval, and Sphinx sequence. Our method demonstrates remarkable adaptability to trajectories it has never seen before.
Figure 11: Estimated trajectories of Blackbird dataset by RoNIN, TLIO, IMO and AirIO (Ours).
A-C Pegasus Dataset
@@
In our experiment, we collected a total of seven trajectories datasets, named Pegasus Dataset. We divided the Pegasus dataset into training and testing sets. Specifically, four trajectories are selected for training and the remaining three trajectories are testing. They are illustrated in 12 and 13. We provide a detailed overview of each trajectory as follows.
@@ A-D EuRoC Dataset
A-E Ablation Study in Pegasus and EuRoC dataset
TABLE V: Ablation study on the EuRoC and Pegasus datasets comparing different feature representations. Evaluation metric: RTE (Unit: ).
@@ A-E Ablation Study in Pegasus and EuRoC dataset
TEST_1 2.783 2.971 2.134 1.694 1.561 1.017
TEST_2 2.704 3.007 1.961 2.339 2.314 1.905
TEST_3 3.274 3.350 1.298 1.969 0.672 0.396
Avg. 2.920 3.109 1.798 2.001 1.516 1.106
A-F Ablation Study on Model Compression
@@
To quantify the performance degradation as the model is compressed, we define the degradation ratio for ATE and RTE. A higher degradation ratio indicates a larger drop in performance. As shown in VI, the model under body frame shows smoother degradation in both ATE and RTE.
TABLE VI: Ablation study on the Blackbird, Pegasus, and EuROC datasets comparing compressibility of models under body frame and global frame. Evaluation metric: ATE (Unit: ), RTE(Unit: ), and Degradation.

### v2

@@ IV-B1 Evaluation on Blackbird Dataset
For the seen sequences, as shown in Table. I, RoNIN and TLIO fail to capture the aggressive drone maneuvers as they are designed for pedestrian navigation with clear slow motion patterns. AirIO significantly outperforms these baselines, improving by 66.1% in ATE and 71.3% in RTE over TLIO.
@@
Furthermore, AirIO exceeds IMO by 56.6% in ATE and 70.2% in RTE relying solely on IMU measurement without integrating external thrust information. We also show that our algorithm can outperform the multi-sensor fusion algorithms like the visual-inertial odometry on the drone racing data. For more details, please check Appendix. A-H.
In addition, it is worth noting that AirIO demonstrates remarkable generalizability to unseen sequences, as shown in Fig. 7. While other methods suffer from significant drift, AirIO consistently maintains precise estimations. The baseline methods tend to overfit the training data, performing well within the training distribution but failing to generalize to UNSEEN data. Compared to IMO, AirIO shows a substantial improvement of 85.5% in ATE and 86.0% in RTE.
IV-B2 Evaluation on EuRoC Dataset
We then evaluate the proposed method on EuRoC dataset, as shown in Table. II. AirIO outperforms both RoNIN and TLIO by 52.9% and 54.4% in ATE, respectively. Thanks to effective uncertainty modeling, AirIO leverages uncertainties learned from both the IMU kinematic model and motion network, ensuring optimal data fusion and improved performance. After integrating EKF, AirIO further improves accuracy, showing an average additional 17.4% improvement in ATE.
@@ IV-B2 Evaluation on EuRoC Dataset
IV-B3 Real time deployment and Analysis
We evaluate the real-time performance of our algorithm on both an NVIDIA Jetson Orin AGX and a desktop with a low-end RTX 2060 GPU, using an IMU operating at 200 Hz. On Orin, AirIO achieves an average runtime of 74.23 ms, while on the desktop, it runs in 28.92 ms. These results indicate the feasibility of our algorithm for real drone deployment.
@@ V Ablation Study
We perform an ablation study to investigate the impact of different feature representations. Experiments with the following input representation scenarios are conducted on the real-world dataset EuRoC and the simulated dataset Pegasus.
@@
    Body: The network takes body-frame IMU measurements defined in (1) as input.
@@
    Global: The network takes global-frame IMU measurements as input. The global-frame IMU is transformed from raw IMU data, as defined in (2).
@@
    Body + Attitude: Building upon i. , the drone orientation is encoded as auxiliary inputs to the network.
@@
    Global + Attitude: Building upon ii. , the drone orientation is encoded as auxiliary inputs to the network.
@@
[…]
@@
[…]
To ensure a fair comparison, scenarios i. , iii. , and v. leverage ground-truth orientation to transform estimated velocity into global frame.
Table. III shows the results of the ablation study for all variants. More results can be found in Appendix. A-F. From these results, we address the following questions:
@@ V-A How Effective is the Body-frame Representation?
Compared to ii. Global, i. Body demonstrates significantly higher precision on the Pegasus. This new input representation results in an average ATE improvement of 74.1%. Scenario i. on EuRoC dataset and Blackbird Unseen sequences exhibit similar trends.
We further evaluate the system performance under two representation using the Accuracy Area Under the Curve (AUC) metric, which quantifies the performance across various error thresholds. Specifically, we calculate the relative position error over a 5-second interval. As shown in Fig. 8, on both datasets, scenario i. achieves higher accuracy across all error thresholds and reaches near-perfect accuracy much more quickly. These results highlight the effectiveness of body-frame representation which efficiently captures implicit attitude information of IMU data. Simply encoding the body frame IMU features improves the system performance.
@@
Seq. Global (2) $-{}^{\mathcal{G}}\mathbf{g}$ Global Body (1) $-\mathbf{R}_{i}^{T}{}^{\mathcal{G}}\mathbf{g}$ Global $+$Attitude Body Body $+$Attitude
EuRoC MH02 18.048 17.892 16.136 9.522 4.537 2.531
@@
Pegasus TEST_1 22.115 9.689 9.534 6.151 5.225 3.418
TEST_2 15.745 15.185 14.729 5.373 4.489 5.719
TEST_3 19.467 26.960 3.580 14.545 3.734 1.786
Avg. 19.109 17.278 9.281 8.690 4.483 3.641
Blackbird Ampersand 21.93 17.863 18.639 14.487 1.977 2.492
Sid 11.866 7.915 5.686 2.307 1.108 0.651
Oval 7.212 6.743 1.656 7.354 1.353 0.913
Sphinx 3.327 3.572 2.117 2.784 1.266 1.163
BentDice 15.5 8.876 6.989 9.626 1.915 1.249
Avg. 11.967 8.994 7.018 7.312 1.524 1.294
V-B How Effective is the Attitude Encoding?
As illustrated in Fig. 9, on the EuRoC dataset, iii. Body + Attitude achieves a further 30.3% improvement over the scenario i. , outperforming scenario ii. by 67.3% in ATE. Additionally, we evaluate attitude encoding on global-frame representation iv. Global + Attitude. This also yields an improvement of approximately 29.6% over scenario ii. on EuRoC dataset, validating the efficiency of encoding attitude information. Encoding the drone’s attitude information enhances the network’s ability for state estimation.
Figure 9: Performance of MH_04_difficult in EuRoC dataset across different representations. ii. Global (Left) is suboptimal. Using IMU data in i. Body (Center) improves significantly. Encoding UAV orientation iii. Body + Attitude (Right) further improves the performance.
V-C Shall We Keep Gravity in the Feature Representation?
Many learning-based IO methods remove gravity from the IMU encoding [26]. However, as discussed in Section. III-A, gravitational acceleration is a critical component for body-frame representation as it explicitly embeds drone rotation through the term $\mathbf{R}^{T}_{i}{}^{\mathcal{G}}\mathbf{g}$, as shown in 1. After removing this rotation-coupled gravity term, scenario i. loses essential attitude information. Table. III demonstrates that v. Body –$\mathbf{R}^{T}_{i}{}^{\mathcal{G}}\mathbf{g}$ causes significant performance degradation: the ATE increases by 90.0% on EuRoC, 107.0% on Pegasus dataset, and 360.5% on Blackbird’s Unseen sequences.
@@ VI Conclusion & Discussion
In this work, we investigate an effective representation of learning-based IO for multirotor UAVs and propose a solution that relies solely on IMU data. We identify that the commonly used global-frame representation is less effective for dynamic, agile maneuvers due to the less observable attitude information compared to IMU data represented in body-frame. Based on this, we develop the AirIO system by explicitly encoding attitude information and integrating it with an EKF. We show that our approach outperforms the state of the art [6] in various scenarios and environments without relying on additional sensors or control information. Our method achieves on average 66.7% improvement in accuracy. Furthermore, explicitly encoding attitude information into the motion network results in an additional 23.8% improvement. We validate the real-time capability of AirIO on an NVIDIA Jetson AGX Orin. Future work includes deploying the system on real UAVs for closed-loop evaluation.
@@ Appendix A Appendix
A-A Real-Time Performance Analysis
To assess the real-time performance of AirIO, we implemented a first-in-first-out (FIFO) IMU buffer to store IMU measurements published at 200 Hz. In our setup, the buffer size is set to 1000, corresponding to approximately 5 seconds of data. At each iteration, we input the entire buffer into the AirIO and AirIMU networks, but only select the outputs corresponding to the newly received (unseen) IMU frames. After network inference, the learned velocity, IMU corrections, and associated covariances are passed to the EKF for fusion. Thus, the total inference time includes both the network computation and the EKF update.
Table. IV shows the comparison of real-time inference time of our approach on different platforms. Overall, AirIO provides the state updates for 28.92 ms on low-end GPU and 74.23 ms on the Orin AGX, which is feasible for real drone deployment. We also notice that most of the inference time is spent on processing the neural network of AirIO, which highly depends on the IMU buffer size.
TABLE IV: Average 1-run inference time (ms) on different platforms for EuRoC.
Platform Host Machine Jetson AGX Orin
CPU Intel Core i9-12900K 12-core ARM Cortex-A78AE
GPU NVIDIA RTX 2060 (6GB) Ampere GPU
OS + Env Ubuntu 22.04, CUDA 12.6 JetPack 3.6
Avg. Inference Time (ms) 28.92 74.23
A-B BlackBird Dataset
@@
A-B1 Seen and Unseen sequences separation
@@
A-B2 Training and Testing Sequences Separation
@@
TABLE V: Separation of trajectory sequences into SEEN and UNSEEN categories, and their respective allocations to training, validation, and testing sets.
@@
A-C Qualitative Evaluation for Blackbird dataset
We present more details on the evaluation of the Blackbird dataset. As shown in 10 and 11, we showcase seven additional trajectories that further highlight our method’s performance. Our method achieves superior performance in the SEEN sequences. For more than half of these sequences, it outperforms existing methods that rely on additional information in addition to IMU measurements, while our method uses only IMU data. When evaluating UNSEEN sequences, our model outperforms all existing methods on all sequences, demonstrating its remarkable adaptability.
Figure 10: SEEN Sequences. From top to bottom: Clover, Egg, Winter and Star. Our method demonstrates robust performance without requiring any additional sensor information.
Figure 11: UNSEEN Sequences. From top to bottom: bentDice, Oval, and Sphinx. Our method demonstrates remarkable adaptability to trajectories it has never seen before.
A-D Pegasus Dataset
@@
In our experiment, we collected a total of seven trajectories datasets, named Pegasus Dataset. We divided the Pegasus dataset into training and testing sets. Specifically, four trajectories are selected for training and the remaining three trajectories are testing. They are illustrated in LABEL:fig:trainmain and 13. We provide a detailed overview of each trajectory as follows.
@@ A-E EuRoC Dataset
A-F Ablation Study in Pegasus and EuRoC dataset
TABLE VI: Ablation study on the EuRoC and Pegasus datasets comparing different feature representations. Evaluation metric: RTE (Unit: ).
@@ A-F Ablation Study in Pegasus and EuRoC dataset
TEST_1 2.783 2.971 2.134 1.694 1.137 1.017
TEST_2 2.704 3.007 1.961 2.339 2.162 1.905
TEST_3 3.274 3.350 1.298 1.969 0.584 0.396
Avg. 2.920 3.109 1.798 2.001 1.294 1.106
Blackbird
Ampersand 8.933 6.845 11.379 8.374 1.254 1.185
Sid 7.768 6.151 3.575 2.72 0.737 0.504
Oval 4.154 3.936 1.094 4.555 0.690 0.558
Sphinx 4.637 3.647 1.695 2.875 0.888 1.021
BentDice 6.998 5.587 5.427 5.578 1.389 0.871
Avg. 6.498 5.233 4.634 4.82 0.992 0.828
A-G Ablation Study on Model Compression
@@
To quantify the performance degradation as the model is compressed, we define the degradation ratio for ATE and RTE. A higher degradation ratio indicates a larger drop in performance. As shown in VII, the model under body frame shows smoother degradation in both ATE and RTE.
TABLE VII: Ablation study on the Blackbird, Pegasus, and EuROC datasets comparing compressibility of models under body frame and global frame. Evaluation metric: ATE (Unit: ), RTE(Unit: ), and Degradation.
@@
A-H Comparison with Visual Inertial Odometry
We have added the results of OpenVINS [32] as a reference. Specifically, we incorporate the quantitative results reported by IMO [6] for the BlackBird dataset and by QVIO [33] for the EuRoC dataset.
As shown in the Table. VIII, on the EuRoC dataset, the VIO system exhibits a significant advantage over IO methods due to the additional vision modality, achieving a 64.1% improvement in ATE. However, on the BlackBird dataset, our method achieves comparable performance across all sequences. Notably, Blackbird dataset features peak velocities up to 8 $\mathrm{m/s}$, leading to motion blur and large optical flows if using vision-based approaches, which make feature tracking challenging and degrade the performance of VIO systems.
TABLE VIII: The ATE (Unit: ) and RTE (Unit: ) on the Blackbird dataset and the EuRoC dataset. Seen are sequences where the training and testing datasets are derived from the same trajectory.
Seq. OpenVINS§ [32] TLIO IMO AirIO Net AirIO EKF
RTE ATE RTE ATE RTE ATE RTE ATE RTE ATE
EuRoC
MH02 - 0.9 2.121 5.902 2.451 7.281 0.936 4.917 0.987 2.478
MH04 - 1.36 4.542 8.586 5.498 8.626 1.093 2.726 1.005 2.308
V103 - 1.68 1.695 3.24 2.58 7.863 1.519 3.844 1.552 3.05
V202 - 1.24 2.303 7.445 2.783 6.26 1.303 4.823 1.313 4.206
V101 - 0.53 1.918 8.576 1.946 4.814 1.137 2.917 1.103 3.844
Avg. - 1.14 2.516 6.75 3.052 6.969 1.198 3.846 1.192 3.177
BlackBird Seen
    Clover - 0.5 0.797 1.464 0.681 0.381 0.368 0.434 0.391 0.367
    Egg - 1.07 2.398 2.227 0.828 1.153 0.391 0.713 0.344 0.408
    halfMoon - 0.37 0.475 0.956 0.24 0.761 0.256 0.491 0.253 0.457
    Star - 2.78 0.784 0.68 3.066 2.13 0.401 0.442 0.341 0.477
    Winter - 0.12 0.755 0.616 0.206 0.219 0.147 0.348 0.164 0.307
Avg. - 0.97 1.042 1.189 1.004 0.929 0.312 0.486 0.299 0.403
§ Visual-Inertial Odometry (VIO) method.